ACTION TO BE DONE:
- Ask Joel the delays for Hositrad
- Ask Joel if the BPM clamps shall maintain fixed or sliding the cables
- Change the screws of the RF collar for serie production (Luca and team)
- Harmonize the tank thread for BPM clamps in M3 for the serie (Luca and team)
- Verify the thickness of the Ag deposit on the TCTPXH screws (Romain)
- Measurement of the CuCD blocks (Carlotta and Dylan)
- Take the BPM connectors at Joel office (Dylan)
- Ask the delays from Plansee (Federico)
- Emilien will sent the tanks metrology reports to Dylan
OTHER ACTIONS:
TCC:
- Julabo technician will arrive 2nd of February
- Protection cover under production
- A test to remove water avec TCC test must be done
BPM flanges:
- Tomography will be performed on the connectors once the connector will be extracted. A connector that is leaking after welding will be analysed
- Extraction ongoing, connector will be removed from the flange on this day
RF fingers:
- Support ongoing, fingers under chemical cut
- Adjust length of the RF fingers testbench
Design:
- TCLP, TCSPM, TCTPM folders: not started
- Post-proto modifications regarding Dylan’s Excel: ongoing for jaws, done for mechanical table, almost finished for feet
- Switch integration: drawing to be done
- TCTPXV review and modifications for serie: Emanuel working on it
- Mechanical table procedure check: ongoing
- Gabarit for cooling pipes production and assembly: ongoing
- Copper wedges for jaws posing during vacuum test (STI)
Assembly:
- TCLPX jaws under metrology to perform flatness measurements post-bakeout
- TCTPXH jaws disassembled, and copper parts came back from metrology -> everything good
- Tank sub-components assembled, waiting for the tanks
- Mechanical tables ready for assembly
- Vireur is ready in clean room
- Propose a date for Gregory and Jan for assembly test of the interconnects
- TCTPXH clamps and screws now ok -> screws Ag + thread were not very good
Procurement:
- Inermet blocks production ongoing, arriving at CERN mid-February
- CuCr1Zr taperings for end of January
Production MME:
- RF fingers ongoing -> new delays from Emilien -> S7: perfect!
- Tank cover arriving mid-March (input from Emilien) then cleaning
- Nanoker will re-machine the wrong blocks
Production STI:
- Jaw flatness measurement tool: waiting for the washer, TCLPX measurement on Tuesday
- Remaining feets components: ongoing
- Transitions and pinch-off: not started
- Lifting ears: not started
Update on planning:
- Collimators shall be operational for July
- TCLPX prevision: September
- TCTPXH prevision: November
- RF fingers, Inermet, TCC test bench are on the critical path
Update for the serie:
- Create a tool to avoid DN300 to pull bellows during tank leak test
- Reduce height and width of the clamps
- Harmonize the tank thread for BPM clamps in M3 for the serie
- Change the screws of the RF collar for serie production
Check to put an elastic element as TCLPX for TCTPXH and avoid overlength of the blocks
There are minutes attached to this event.
Show them.